# Expressive Temporal Specifications for Reward Monitoring

### v1

@@ 2 Background
Reinforcement Learning (RL) environments are typically modelled as Markov decision processes (MDPs), defined as tuples $\mathcal{M}=(\mathcal{S},A,\Pr,\mathcal{R},\gamma)$, where $\mathcal{S}$ is a set of states; $A$ is a set of actions; $\Pr:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition kernel, with $\Delta(\mathcal{S})$ being the set of probability distributions over $\mathcal{S}$, $\mathcal{R}:\mathcal{S}\times A\times\mathcal{S}\to\mathbb{R}$ is the (Markovian) reward function, and $\gamma\in[0,1]$ is the discount factor.
@@ 3.3 Monitor Synthesis
Let $\varphi$ be an ltlf[][$\mathcal{F}$] formula, $\mathcal{A}_{\varphi}$ the QRM constructed from $\varphi$ as per Alg. 1, and $\lambda$ a finite trace of length $n$. Let $\mathrm{val}_{\mathcal{A}_{\phi}}(\,\varphi,\,i\,)$ denote the value stored in the reward register at time index $i$. Then, for each index $1\leq i\leq n$,
@@ 4 Experimental Evaluation
with span $N$. Define checkpoints every $N$ steps: $C_{i}\stackrel{{\scriptstyle def}}{{=}}E_{iN}$.
@@ B.4 Proof of Correctness
    $\varphi=p$, for an atomic proposition, the semantic value at a given time step $i$ is $\bigl[\![\,p,i\,]\!\bigr](\lambda)$. The monitor constructed transitions from the initial state to an empty self-looping state with no further register updates. The initial state provides a reward output of the current valuation of $p$ from the labelling function $\mathcal{L}_{p}(s)$ of the MDP, and this is stored permanently in the reward register, which corresponds to evaluating a proposition $p$ at index $i$.
@@
    Conjunction $\varphi=\psi_{1}\land\psi_{2}$, by constructing the respective submonitors of $\mathcal{A}^{\psi}1$ and $\mathcal{A}^{\psi}2$, taking the Cartesian product of the state and transition space, tracking the propositional values of each formula, we append an additional register that tracks the minimum between the two formulae such that $\min(\psi_{1},\psi_{2})$ and use this as the reward register, which provides the correct output for time step $i$.
@@
    Next $\varphi=X\psi$ is almost identical to the base case of an atomic proposition, but requires accessing the next timestep $i+1$ instead of $i$. Whilst requiring a future time step’s value is typically an issue with runtime verification, when performing an action we have the triple ($s,a,s^{\prime}$), therefore the labelling of $s^{\prime}$ is considered for $i+1$.
@@ Appendix C Environment Specification Details
In all experiments, we used an exponentially moving average convergence test (see Definition 8) with span $N=32$ episodes, forming checkpoints every $N$ episodes and declaring convergence once the last $P=5$ checkpoint intervals showed less than a 1 relative improvement in the EMA of returns. The relative tolerance was automatically adapted to the recent noise in the checkpoint sequence, clipped to the range $\rho_{\min},\rho_{\max}]=[0.002,0.02]$, and we did not enforce a minimum EMA level or exclude negative returns.

### v2

@@ 1 Introduction
footnotetext: * This paper is the extended version (with appendix) of the paper in the Proceedings of the 40th AAAI Conference on Artificial Intelligence (AAAI-26).
@@ 2 Background
Reinforcement Learning (RL) environments are typically modelled as Markov decision processes (MDPs), defined as tuples $\mathcal{M}=(\mathcal{S},A,\Pr,\mathcal{R},\gamma)$, where $\mathcal{S}$ is a set of states; AA is a set of actions; $\Pr:\mathcal{S}\times A\to\Delta(\mathcal{S})$ is the transition kernel, with $\Delta(\mathcal{S})$ being the set of probability distributions over $\mathcal{S}$, $\mathcal{R}:\mathcal{S}\times A\times\mathcal{S}\to\mathbb{R}$ is the (Markovian) reward function, and $\gamma\in[0,1]$ is the discount factor.
@@ 3.3 Monitor Synthesis
Let $\varphi$ be an ltlf[][$\mathcal{F}$] formula, $\mathcal{A}_{\varphi}$ the QRM constructed from $\varphi$ as per Alg. 1, and $\lambda$ a finite trace of length $n$. Let $\mathrm{val}_{\mathcal{A}_{\phi}}(\,\varphi,\,i\,)$ denote the value stored in the reward register at time index ii. Then, for each index $1\leq i\leq n$,
@@ 4 Experimental Evaluation
with span NN. Define checkpoints every NN steps: $C_{i}\stackrel{{\scriptstyle def}}{{=}}E_{iN}$.
@@ B.4 Proof of Correctness
    $\varphi=p$, for an atomic proposition, the semantic value at a given time step $i$ is $\bigl[\![\,p,i\,]\!\bigr](\lambda)$. The monitor constructed transitions from the initial state to an empty self-looping state with no further register updates. The initial state provides a reward output of the current valuation of $p$ from the labelling function $\mathcal{L}_{p}(s)$ of the MDP, and this is stored permanently in the reward register, which corresponds to evaluating a proposition $p$ at index ii.
@@
    Conjunction $\varphi=\psi_{1}\land\psi_{2}$, by constructing the respective submonitors of $\mathcal{A}^{\psi}1$ and $\mathcal{A}^{\psi}2$, taking the Cartesian product of the state and transition space, tracking the propositional values of each formula, we append an additional register that tracks the minimum between the two formulae such that $\min(\psi_{1},\psi_{2})$ and use this as the reward register, which provides the correct output for time step ii.
@@
    Next $\varphi=X\psi$ is almost identical to the base case of an atomic proposition, but requires accessing the next timestep $i+1$ instead of ii. Whilst requiring a future time step’s value is typically an issue with runtime verification, when performing an action we have the triple ($s,a,s^{\prime}$), therefore the labelling of $s^{\prime}$ is considered for $i+1$.
@@ Appendix C Environment Specification Details
In all experiments, we used an exponentially moving average convergence test (see Definition 8) with span $N=32$ episodes, forming checkpoints every NN episodes and declaring convergence once the last $P=5$ checkpoint intervals showed less than a 1 relative improvement in the EMA of returns. The relative tolerance was automatically adapted to the recent noise in the checkpoint sequence, clipped to the range $\rho_{\min},\rho_{\max}]=[0.002,0.02]$, and we did not enforce a minimum EMA level or exclude negative returns.